OpenStudy (anonymous):

Someone please check my answer Consider this function: f(x)=(x^2-9)(4x^2+4x-3) What are the zeros of the function? A)-3, -2/3, 2, 3 B)-3, -2/3, -2, 3 C)-3, -3/2, 1/2, 3 D)-3, -1/2, 3/2, 3 i believe is is C but i dont think its right... very confused please help will give medals!!

jimthompson5910 (jim_thompson5910):

To find the zeros, or roots, of the function you replace f(x) with 0 and solve for x f(x)=(x^2-9)(4x^2+4x-3) 0=(x^2-9)(4x^2+4x-3) (x^2-9)(4x^2+4x-3) = 0 x^2-9 = 0 or 4x^2+4x-3 = 0 From here you have to solve x^2-9 = 0 and you also have to solve 4x^2+4x-3 = 0 You can use the quadratic formula for each

OpenStudy (anonymous):

Hmmmmm..... I do not know about this one... i will do a bit of researching and get back to you maybe @jim_thompson5910 can help?

OpenStudy (anonymous):

how would I use the quadratic equation when there isn't an a b and c?? I'm sorry if I sound stupid

jimthompson5910 (jim_thompson5910):

x^2-9 = 0 is the same as 1x^2+0x+(-9) = 0 so here a = 1 b = 0 c = -9

OpenStudy (anonymous):

oh yea thats right. so if i use the quadratic equation i will get the answer?

jimthompson5910 (jim_thompson5910):